Darfangel:

can someone help me pls :0

Darfangel:

so i dont understand this

Darfangel:

1 attachment
kristen00001:

it might be 10/3 i really dont know but thats my guess

fireballfireball:

oh man sorry bro im only in 8th grade i would help tho

CripQUEZZ:

do yk how to find slope with two ordered pairs?

urmum:

its y2-y1/x2-x1

JNJackson02:

im in 8th grade and we starting on this

fireballfireball:

not yet for me

kristen00001:

im 9th grade and im doing this

CripQUEZZ:

the formula is y2-y1/x2-x1

Darfangel:

okie

CripQUEZZ:

so 5 + 15 / -2 + 8

CripQUEZZ:

that will give u ur answer

Darfangel:

@cripquezz wrote:
so 5 + 15 / -2 + 8
okie thx

urmum:

@cripquezz wrote:
so 5 + 15 / -2 + 8
yep you do the math

CripQUEZZ:

@darfangel wrote:
@cripquezz wrote:
so 5 + 15 / -2 + 8
okie thx
no problem

JNJackson02:

5- -15=-10 10/6=5/3 so, 5/3 is the answer -2-8=8

CripQUEZZ:

@jnjackson02 wrote:
5- -15=-10 10/6=5/3 so, 5/3 is the answer -2-8=8
no u have to do keep change change on the top keep the 5 and change the wo negatives

CripQUEZZ:

two*

Darfangel:

so what me confued??

CripQUEZZ:

do it how i showed u

CripQUEZZ:

5 + 15 / -2 + 8

CripQUEZZ:

that will give u ur slope

urmum:

@cripquezz wrote:
5 + 15 / -2 + 8
you subract, not add

CripQUEZZ:

@urmum wrote:
@cripquezz wrote:
5 + 15 / -2 + 8
you subract, not add
u have to add bc they are negative numbers

Darfangel:

is the slash a fraction?

CripQUEZZ:

the formula is y2 - y1 / x2 - x1 if the numbers are negative u have to do keep change change and change it to a positive

CripQUEZZ:

@darfangel wrote:
is the slash a fraction?
yes it is

Darfangel:

so 5/3 would be the answer?

CripQUEZZ:

yes

Darfangel:

oof it said the answer was 10/3 thx for the help tho!

kristen00001:

told ya

JNJackson02:

lol

kristen00001:

ppl never listen

CripQUEZZ:

was it supposed to be in simplist form?

CripQUEZZ:

oh it didnt say to put it in simplist form lol

Darfangel:

yup i think

CripQUEZZ:

i think it was jus supposed to be the answer unsimplified

CripQUEZZ:

sorry lol

kristen00001:

you shoulve listen to me in the 1st place and this convo would be done with

CripQUEZZ:

i doesnt matter lmao technically theyre still right its jus simplified